OpenStudy (anonymous):

ABCD is a parallelogram. Its diagonal, AC, is 18 inches long and forms a 20° angle with the base of the parallelogram. Angle ABC is 130°. What is the length of the parallelogram’s base, AB? 8.0 in 11.7 in 27.6 in 40.3 in

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@rhr12

OpenStudy (anonymous):

@wolf1728

OpenStudy (rhr12):

do you know the area of the parallelogram?

OpenStudy (wolf1728):

We can say that the half angle located at the upper right would be 30 degrees. We have 2 angles and an included side so we can use the Law of Sines

OpenStudy (rhr12):

we can eliminate 40 and 26. the confusion is 8 or 11.7.

OpenStudy (wolf1728):

I drew a new diagram

OpenStudy (anonymous):

11.7?

OpenStudy (rhr12):

11.7 i guess

OpenStudy (anonymous):

thats ehat i had first but i wasnt ture so i asked yall

OpenStudy (rhr12):

is that wrong? I had never did something like that. MAybe @wolf1728 will take you to the end

OpenStudy (wolf1728):

Angle B is 130 degrees side b is 18 inches Angle C is 30 degrees side C = (sin(C) * b)/sin B side C = .5 * 18 / 0.76604 side C = 11.7487337476

OpenStudy (wolf1728):

Well I guess it is 11.7 and you've got the proof right there! :-)